# A Gaussian Sliding Windows Regression Model for Hydrological Inference

### v3

@@ 2.5 Autocorrelated residuals
$y_{t}=\sum_{i=1}^{k}\beta_{i}\cdot\left({\bm{x}}_{[t]}\ast\bm{\kappa}^{(i)}\right)+\varphi\varepsilon_{t-1}+\eta_{t},$
leading to
$y_{t}-\varphi y_{t-1}=\sum_{i=1}^{k}\beta_{i}\cdot\left(\left({\bm{x}}_{[t]}-\varphi{\bm{x}}_{[t]}\right)\ast\bm{\kappa}^{(i)}\right)+\eta_{t}.$ (7)
Here, we are exploiting the linearity of the convolution operator. Hence, given $\varphi$, the model can be written as an ordinary least squares model with uncorrelated errors $\eta_{t}$:
$\tilde{y}_{t}=\sum_{i=1}^{k}\beta_{i}\cdot\left(\tilde{x}_{[t]}\ast\bm{\kappa}^{(i)}\right)+\eta_{t},$
@@
  • estimate model parameters $\bm{\beta},\bm{\delta},\bm{\sigma}$ from the original model using $x_{t}$ and $y_{t}$ (i.e., set $\varphi=0$),
@@
  • apply the transformation $\tilde{z}_{t}=z_{t}-\varphi z_{t-1}$ to both $x_{t}$ and $y_{t}$ and obtain final estimates of the model parameters $\bm{\beta},\bm{\delta},\bm{\sigma}$.
@@ 2.6 Uncertainty quantification
$\mathcal{I}(\bm{\beta},\bm{\delta},\bm{\sigma})=\frac{\partial^{2}}{\partial(\bm{\beta},\bm{\delta},\bm{\sigma})^{2}}\left(-\ln L(\bm{\beta},\bm{\delta},\bm{\sigma};\bm{x},\bm{y})\right)$
denotes the observed information. The Hessian can be obtained numerically via the R package numDeriv (Gilbert and Varadhan, 2019). The diagonal elements of $\mathcal{I}^{-1}$ lead to approximate standard errors, which provide the uncertainty quantification reported below for the two watersheds studied. These watersheds have very long time series to support the asymptotic argument.
@@ 3.1 Experimental Setup
We employ 15 distinct setups (five 1-, 2-, and 3-window setups, respectively). A higher number of windows is not common for hydrological models due to the limited number of possible flow paths. Each setup uses a set of parameters independently sampled from $\delta_{i}\sim U(0,20)$, and $\sigma_{i},\beta_{i}\sim U(1,5)$ for any $i\in\{1,\dots,k\}$, where $U(a,b)$ denotes a uniform distribution on the interval $[a,b]$. Each 1-, 2- and 3-window setup is repeated at 5 distinct levels of measurement noise $\alpha\in\{0.05,0.25,0.5,0.75,0.95\}$ to control the error relative to the signal, for a total of 75 simulation experiments.
@@ 3.2 Evaluation Metrics
    $\sum\limits_{i=1}^{k}\beta_{i}\bm{\kappa}_{i},$
@@ 3.3 Results for uncorrelated errors
Fig. 6 compares the weights at discrete time points of the ground truth and estimated kernels for the middle noise level $\alpha=0.5$ in each setup. Dotted vertical lines indicate the true or estimated window centres $\delta_{i}$ along the time axis. In all 1-, 2-, and 3-window setups, window positions and sizes were accurately predicted. Moreover, dominant peaks and general shapes are reconstructed in all scenarios. Since we allow non-integer numbers for delta, the discretized weights are not always symmetric w.r.t. the measured time points. Overall, the combined kernel matches with the ground truth very accurately, and the general separation of the windows is acceptable.
@@ 3.4 Results for autocorrelated errors
$=\dots=\sum\limits_{i=0}^{t-1}\varphi^{2i}\text{Var}(\eta_{t-i})$
@@ 4 Real-World Data Experiments
In the second part of our experiments, we evaluate the performance of the Gaussian SWR model on two real-world examples. Both watersheds are located on the west coast of North America, with the first catchment representing the the Koksilah River which is located in Cowichan, British Columbia, Canada. The second catchment represents Big Sur River in central California, United States. The available dataset consists of 39 hydrological years for both watersheds. The Koksilah River has been investigated in previous works (Janssen et al., 2023).
@@ 4.2 Results
watershed no window no $\beta_{i}$ $\delta_{i}$ $\sigma_{i}$
@@
The Gaussian SWR model selects $k=3$ windows for both Koksilah and Big Sur Rivers. Fig. 9(a) shows a clear dominance of the short-term history (a window close to a lag of 1 day) for the Koksilah River. Paired with the fact that only three highly overlapping windows were needed to model the catchment with high accuracy (Table 5), we can conclude that the Koksilah River is driven by simple hydrological processes. The third window is very wide with lower weight, and the first two windows have $\delta\simeq 1$, thus the combined kernel for watershed 1 resembles a truncated Student-$t$ distribution: the combined kernel remains unimodal, but the probability mass is more dispersed over the long-term lags compared to a Gaussian kernel. Moreover, a low lag value suggests a relatively simple catchment, which aligns with the observation that the Koksilah River is situated in a region characterized by extremely high precipitation and minimal aridity (Janssen and Ameli, 2021). From previous works we can hypothesize and now confirm that wet, non arid, catchments tend to be extremely simple with fast response times (Schoppa et al., 2020; Spieler et al., 2020). Since the bulk of precipitation transitions into streamflow in about 1 day, this catchment is likely dominated by shallow subsurface flow. Large amounts of precipitation in this catchment fully saturate and connect the shallow soil layers for much of the year, resulting in any excess precipitation pushing the water already in the soils towards the stream within one day.
@@ 5 Discussion and Conclusion
Relative to the more general DLM, the predictive performance of the Gaussian SWR model is comparable on the datasets used in our real-world experiments. When varying the number of time lags $q$, the DLM implementation by Demirhan (2020) achieves test-set $R^{2}$ scores of 0.7 ($q=5$), 0.72 ($q=10$ and $q=20$) for watershed 1 and 0.51 ($q=5$), 0.55 ($q=10$), 0.58 ($q=20$) for watershed 2. Thus, the parametric assumption taken by Gaussian SWR, i.e. restricting the kernel shape to a mixture of Gaussians, does not significantly affect the predictive performance. However, the favourable parametrization employed by the Gaussian SWR model allows for straightforward interpretations of the model parameters: the window position and shape parameters are directly transferable to the concept of distinct flow paths in hydrology and deliver information on the peak lag of the runoff and its celerity distribution, respectively. Regression parameters $\bm{\beta}$ give indications of the relative importance of each flow path: while a dominant window with short lag typically models a strong short-term (overland) flow, well-distinguished windows with longer time lags indicate the presence of distinct long-term effects such as sub-surface or groundwater flows.

### v4

@@ 2.5 Autocorrelated residuals
$y_{t}$ $=\sum_{i=1}^{k}\beta^{(i)}\cdot\left({\bm{x}}_{[t]}\ast\bm{\kappa}^{(i)}\right)+\varphi\varepsilon_{t-1}+\eta_{t},\text{and}$
$y_{t-1}$ $=\sum_{i=1}^{k}\beta^{(i)}\cdot\left({\bm{x}}_{[t-1]}\ast\bm{\kappa}^{(i)}\right)+\varepsilon_{t-1}$
In order to convert this setup into a model with Gaussian white noise errors, we investigate the expression
$y_{t}-\varphi y_{t-1}$ $=\sum_{i=1}^{k}\beta^{(i)}\cdot\left(\bm{x}_{[t]}\ast\bm{\kappa}^{(i)}-\varphi\bm{x}_{[t-1]}\ast\bm{\kappa}^{(i)}\right)+\eta_{t}$
$=\sum_{i=1}^{k}\beta^{(i)}\cdot\left(\sum\limits_{s=1}^{t}\left(x_{s}-\varphi x_{s-1}\right)\kappa^{(i)}_{t-s}\right)+\eta_{t}.$
The term $x_{0}\cdot\kappa^{(i)}_{t}$ originating from index $s=0$ is ignored since $\bm{\kappa}^{(i)}$ is padded with 00 at the end (when being adjusted to the length of $\bm{x}_{[t]}$). Hence, given $\varphi$, the model can be written as an ordinary least squares model with uncorrelated errors $\eta_{t}$:
$\tilde{y}_{t}=\sum_{i=1}^{k}\beta^{(i)}\cdot\left(\tilde{\bm{x}}_{[t]}\ast\bm{\kappa}^{(i)}\right)+\eta_{t},$
@@
  • estimate model parameters $\bm{\beta},\bm{\delta},\bm{\sigma}$ from the original model using $\bm{x}_{[t]}$ and $\bm{y}_{[t]}$ (i.e., set $\varphi=0$),
@@
  • apply the transformation $\tilde{z}_{t}=z_{t}-\varphi z_{t-1}$ to both $\bm{x}_{T}$ and $\bm{y}_{T}$ and obtain final estimates of the model parameters $\bm{\beta},\bm{\delta},\bm{\sigma}$.
@@ 2.6 Uncertainty quantification
$\mathcal{I}(\bm{\beta},\bm{\delta},\bm{\sigma})=\frac{\partial^{2}}{\partial(\bm{\beta},\bm{\delta},\bm{\sigma})^{2}}\left(-\ln L(\bm{\beta},\bm{\delta},\bm{\sigma};\bm{x}_{T},\bm{y}_{T})\right)$
denotes the observed information, i.e., the negative second derivative of the log-likelihood evaluated at the MLE parameter values. The Hessian can be obtained numerically via the R package numDeriv (Gilbert and Varadhan, 2019). The diagonal elements of $\mathcal{I}^{-1}$ lead to approximate standard errors, which provide the uncertainty quantification reported below for the two watersheds studied. These watersheds have very long time series to support the asymptotic argument.
@@ 3.1 Experimental Setup
We employ 15 distinct setups (five 1-, 2-, and 3-window setups, respectively). A higher number of windows is not common for hydrological models due to the limited number of possible flow paths. Each setup uses a set of parameters independently sampled from $\delta^{(i)}\sim U(0,20)$, and $\sigma^{(i)},\beta^{i}\sim U(1,5)$ for any $i\in\{1,\dots,k\}$, where $U(a,b)$ denotes a uniform distribution on the interval $[a,b]$. Each 1-, 2- and 3-window setup is repeated at 5 distinct levels of measurement noise $\alpha\in\{0.05,0.25,0.5,0.75,0.95\}$ to control the error relative to the signal, for a total of 75 simulation experiments.
@@ 3.2 Evaluation Metrics
    $\sum\limits_{i=1}^{k}\beta^{i}\bm{\kappa}^{(i)},$
@@ 3.3 Results for uncorrelated errors
Fig. 6 compares the weights at discrete time points of the ground truth and estimated kernels for the middle noise level $\alpha=0.5$ in each setup. Dotted vertical lines indicate the true or estimated window centres $\delta^{(i)}$ along the time axis. In all 1-, 2-, and 3-window setups, window positions and sizes were accurately predicted. Moreover, dominant peaks and general shapes are reconstructed in all scenarios. Since we allow non-integer numbers for delta, the discretized weights are not always symmetric w.r.t. the measured time points. Overall, the combined kernel matches with the ground truth very accurately, and the general separation of the windows is acceptable.
@@ 3.4 Results for autocorrelated errors
$=\sum\limits_{i=0}^{t-1}\varphi^{2i}\text{Var}(\eta_{t-i})$
@@ 4 Real-World Data Experiments
In the second part of our experiments, we evaluate the performance of the Gaussian SWR model on two real-world examples. Both watersheds are located on the west coast of North America, with the first catchment representing the the Koksilah River which is located in Cowichan, British Columbia, Canada. The second catchment represents Big Sur River in central California, United States. The available dataset consists of 39 hydrological years for both watersheds. The Koksilah River has been investigated in previous works (Janssen et al., 2023). Both datasets are available on the aforementioned GitHub repository (entitled sampleWatershed and sampleWatershed2, respectively).
@@ 4.2 Results
watershed no window no $\beta^{(i)}$ $\delta^{(i)}$ $\sigma^{(i)}$
@@
The Gaussian SWR model selects $k=3$ windows for both Koksilah and Big Sur Rivers. Fig. 9(a) shows a clear dominance of the short-term history (a window close to a lag of 1 day) for the Koksilah River. Paired with the fact that only three highly overlapping windows were needed to model the catchment with high accuracy (Table 5), we can conclude that the Koksilah River is driven by simple hydrological processes. The third window is very wide with lower weight, and the first two windows have $\delta\simeq 1$, thus the combined kernel for watershed 1 resembles a truncated Student-$t$ distribution: the combined kernel remains unimodal, but the probability mass is more dispersed over the long-term lags compared to a Gaussian kernel. Moreover, a low lag value suggests a relatively simple catchment, which aligns with the observation that the Koksilah River is situated in a region characterized by extremely high precipitation and minimal aridity (Janssen and Ameli, 2021). From previous works we can hypothesize that wet, non arid, catchments tend to be extremely simple with fast response times (Schoppa et al., 2020; Spieler et al., 2020), which is strengthened by our experimental results. Since the bulk of precipitation transitions into streamflow in about 1 day, this catchment is likely dominated by shallow subsurface flow. Large amounts of precipitation in this catchment fully saturate and connect the shallow soil layers for much of the year, resulting in any excess precipitation pushing the water already in the soils towards the stream within one day.
@@ 5 Discussion and Conclusion
Relative to the more general DLM, the predictive performance of the Gaussian SWR model is comparable on the datasets used in our real-world experiments. When varying the number of time lags $q$, the DLM implementation by Demirhan (2020) achieves test-set $R^{2}$ scores of 0.7 ($q=5$), 0.72 ($q=10$ and $q=20$) for watershed 1 and 0.51 ($q=5$), 0.55 ($q=10$), 0.58 ($q=20$) for watershed 2. Thus, the parametric assumption taken by Gaussian SWR, i.e. restricting the kernel shape to a mixture of Gaussians, does not significantly affect the predictive performance. However, the favourable parametrization employed by the Gaussian SWR model allows for straightforward interpretations of the model parameters: the window position and shape parameters are directly transferable to the concept of distinct flow paths in hydrology and deliver information on the peak lag of the runoff and its celerity distribution, respectively. Regression parameters $\bm{\beta}$ give indications of the relative importance of each flow path: while a dominant window with short lag typically models a strong short-term (overland) flow, well-distinguished windows with longer time lags indicate the presence of distinct long-term effects such as sub-surface or groundwater flows. From a statistical perspective, kernels representing long-term effects may approximate a constant line, thus adding an intercept could be a valid option for the proposed model — for the hydrological interpretation, however, a large (slow-flow) window acts as a more accurate parametrization than a model with intercept.